OpenStudy (lindsey11x16):

how do i factor tan^2x - tan^2x sin ^2x :(

OpenStudy (anonymous):

@AriPotta

OpenStudy (lindsey11x16):

?

jimthompson5910 (jim_thompson5910):

tan^2 is the common factor, so factor that out first tan^2x - tan^2x sin ^2x tan^2x(1 - sin ^2x) Now replace 1-sin^2 with cos^2 (this is one of the many trig identities) to get this answer tan^2x * cos^2x since tan = sin/cos, we can say this tan^2x * cos^2x turns into (sin^2x/cos^2x) * cos^2x but the cosine squared terms will cancel leaving you with just sin^2x

jimthompson5910 (jim_thompson5910):

so in the end, the whole expression is just sin^2x

OpenStudy (lindsey11x16):

where did the 1 come from in this tan^2x(1 - sin ^2x) at the beginning

jimthompson5910 (jim_thompson5910):

i factored that out, it's like saying a - ab is the same as a(1 - b) after you factor out 'a'

jimthompson5910 (jim_thompson5910):

in this case, a = tan^2x and b = sin^2x